OpenStudy (anonymous):

AP Calculus: Slope Fields/ Separate Variables Find the equation of the curve, given the derivative and the indicated point on the curve. dy/dx = -1/x^2 dx

OpenStudy (anonymous):

given point on graph is (1,3)

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Is that \[\frac{dy}{dx}=-\frac{1}{x^2}?\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[dy=-\frac{dx}{x^2}\\ \int dy=-\int \frac{dx}{x^2}\]

OpenStudy (anonymous):

so how would i integrate that

OpenStudy (anonymous):

Well, you can recognize that \(\displaystyle-\frac{1}{x^2}\) is the derivative of \(\displaystyle\frac{1}{x},\) or you can write the RHS as \[-\int x^{-2}dx,\] then apply the power rule for integrating. Recall that the power rule is \[\int x^ndx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

ok so the answer is \[y=\frac{ -1 }{ x^2 } + 3\]

OpenStudy (anonymous):

Not quite. After integrating, you get \[y=\frac{1}{x}+C\]

OpenStudy (anonymous):

Because the derivative of this (in differential form) is \[dy=-\frac{1}{x^2}dx+0\\ dy=-\frac{1}{x^2}dx\]

OpenStudy (anonymous):

so that's the answer? really? i'm so confused :S

OpenStudy (anonymous):

Which part are you confused about?

OpenStudy (anonymous):

that the answer isn't \[y=\frac{ -1 }{ x^2 } + 3\]

OpenStudy (anonymous):

the given point on graph is (1,3)

OpenStudy (anonymous):

I'll go through the solution-finding step-by-step: You're given \[\frac{dy}{dx}=-\frac{1}{x^2}\] Move the dx to the right: \[dy=-\frac{1}{x^2}dx\] Integrate both sides: \[\int dy=\int\left(-\frac{1}{x^2}\right)dx\] I'll use the power rule here; I rewrite 1/x² as x^(-2): \[\int dy=\int\left(-x^{-2}\right)dx\\ y=-\left(\frac{x^{-2+1}}{-2+1}\right)+C\\ y=-\left(\frac{x^{-1}}{-1}\right)+C\\ y=-\left(-\frac{1}{x}\right)+C\\ \color{red}{y=\frac{1}{x}+C}\] Given the point (1,3), you have \[3=\frac{1}{1}+C\\ 3=1+C\\ 2=C\] So, the solution is \[y=\frac{1}{x}+2\] The red is where the misunderstanding seems to take place.

OpenStudy (anonymous):

yeah that's what i put i think i forgot to subtract the 1 to the other side lol

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

ok ty :) can you help me with another one please? ^^

OpenStudy (anonymous):

yep